# revtex4-2Repair the float

### v2

@@ Appendix D Vortex in a gyrotropic superconductor
Now, according to Eq. (11) in the main body, a gradient in the Green’s function is coupled to the spin via the term in the torque $\gamma_{kj}\sigma_{j}\partial_{k}g_{s}$, where $g_{s}$ is the Green’s function in the absence of a spin-galvanic effect. We may distinguish two different kinds of spatial dependence, spatial dependence of the phase and spatial dependence of the magnitude. Indeed, for a general system, we may write
$g_{s}(\bm{r})=e^{-i\phi(\bm{r})\tau_{3}/2}(g_{s0}(\bm{r})\tau_{3}+f_{s0}(\bm{r})\tau_{2})e^{i\phi(\bm{r})\tau_{3}/2},$ (60)
@@
$\bm{h}_{\text{eff},j}=n\bm{\hat{\phi}}_{k}\gamma_{kj}.$ (61)
To derive the symmetry of the spin for generic forms of the spin-galvanic tensor, it is instructive to consider three cases separately, in all of which the tensor has two non-vanishing elements. First of all, we consider $\gamma_{xy}=-\gamma_{yx}$. A well-known example of spin-orbit coupling that leads to a spin-galvanic tensor of this form is Rashba spin orbit coupling. Generally it is the only allowed symmetry in the $C_{3v},C_{4v}$ and $C_{6v}$ point groups. For this symmetry the spin is always perpendicular to the direction of the current and hence around a vortex the spin points in the radial direction. The sign of $n\gamma_{xy}$ determines whether this is radially outward ($>0$) or inward ($<0$). Moreover, the magnitude of the excess spin is homogeneous at a fixed distance from the vortex core, it is independent of $\phi$.
@@
We now compute explicitly the current and excess spin densities by solving the Usadel equation, Eq. (7), assuming an isotropic spin-relaxation $\Gamma_{jk}=\frac{D}{l_{\text{so}^{2}}}\delta_{jk}$, where $l_{\text{so}}$ is the spin-relaxation length, for simplicity of notation. For this we use the Riccati parameterization (we use the notation $a,b$ instead of the usual $\gamma,\tilde{\gamma}$ to avoid confusion with the tensor of the gyrotropic term)
$g=\begin{bmatrix}(1+ab)^{-1}(1-ab)\text{sign}(\omega_{n})&2(1+ab)^{-1}a\\ 2(1+ba)^{-1}b&-(1+ba)^{-1}(1-ba)\text{sign}(\omega_{n})\end{bmatrix}\;.$ (62)
This parametrization guarantees the normalization condition $g^{2}=1$. For a single vortex, in the absence of a gyrotropic tensor the equation for $a$ can be written as (cf. Eq. (7))
$\nabla^{2}a-a\nabla a\cdot\nabla b=2\frac{|\omega|}{D}a+\frac{2\Delta}{D}e^{in\phi}(1-a^{2}-(be^{in\phi}-ae^{-in\phi}))+\frac{1}{l_{\text{so}}^{2}}(a-\sigma_{k}a\sigma_{k}),$ (63)
@@
$\frac{\partial^{2}a}{\partial r^{2}}+\frac{1}{r}\frac{\partial a}{\partial r}+\frac{1}{r^{2}}\frac{\partial^{2}a}{\partial\phi^{2}}-a\Big{(}\frac{\partial a}{\partial r}\frac{\partial b}{\partial r}+\frac{1}{r^{2}}\frac{\partial a}{\partial\phi}\frac{\partial b}{\partial\phi}\Big{)}=\frac{2|\omega|}{D}a-\frac{\Delta}{D}(1-a^{2}-(be^{in\phi}-ae^{-in\phi}))+\frac{1}{l_{\text{so}}^{2}}(a-\sigma_{k}a\sigma_{k})$ (64)
@@
$a_{0}(r,\phi)=a_{s}(r)e^{in\phi},$ (65)
$b_{0}(r,\phi)=a_{s}(r)e^{-in\phi}.$ (66)
The equation for $a_{s}(r)$ is a second order non-linear equation with non-constant coefficients in one variable;
$\frac{\partial^{2}a_{s}}{\partial r^{2}}+\frac{1}{r}\frac{\partial a_{s}}{\partial r}-\frac{n^{2}}{r^{2}}a_{s}-a_{s}\Big{(}\left(\frac{\partial a_{s}}{\partial r}\right)^{2}+\frac{n^{2}}{r^{2}}a_{s}^{2}\Big{)}=\frac{2|\omega|}{D}a_{s}-\frac{\Delta}{D}(1-a_{s}^{2}).$ (67)
@@
$I(r)=\frac{-i}{1+a_{0}b_{0}}(a_{0}\nabla b_{0}-b_{0}\nabla a_{0})=\frac{1}{r}\frac{na_{s}^{2}}{1+a_{s}^{2}}\bm{\hat{\phi}}.$ (68)
We note that while there exists a factor $\frac{1}{r}$ in the expression, the current density does not diverge near the center, since $a_{s}$ has to be at least first order in $r$ following Eq. (67). Hence, the current density vanishes at $r=0$.
@@
$=\frac{2|\omega|}{D}\delta a_{t}+\frac{2\Delta}{D}a_{s}\delta b_{t}+\frac{\Delta}{D}(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi})+\gamma_{o}(\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}-\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{in\phi}\;,$ (69)
@@
$=\frac{2|\omega|}{D}\delta b_{t}+\frac{\Delta}{D}(2a_{s}\delta a_{t}-(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi}))-\gamma_{o}(-\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}+\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{-in\phi}\;.$ (70)
@@
Again exploiting the symmetry of the problem we may write $\delta a_{t}(r,\phi)=(a_{x}(r)\sigma_{x}+a_{y}(r)\sigma_{y})e^{i\phi\sigma_{z}}e^{in\phi}$ and $\delta b_{t}(r,\phi)=\sigma_{y}\delta a_{t}(r,\phi)^{*}\sigma_{y}=-(a_{x}(r)^{*}\sigma_{x}+a_{y}(r)^{*}\sigma_{y})e^{i\phi\sigma_{z}}e^{-in\phi}$ to obtain two coupled non-linear second order differential equations. We first consider the case in which the tensor of the gyrotropic term has the same symmetry as for Rashba spin orbit coupling, i.e. $\gamma_{xy}=-\gamma_{yx}=\gamma_{o}$, while all other terms of the tensor vanish. In this case
@@
$=\frac{2|\omega|}{D}a_{x}+\frac{1}{l_{\text{so}}^{2}}a_{x}+\frac{2\Delta}{D}a_{s}a_{x}+\frac{2\Delta}{D}\text{Re}a_{x}+\frac{in}{r}\gamma_{o}a_{s}\;,$ (71)
@@
$=\frac{2|\omega|}{D}a_{y}+\frac{1}{l_{\text{so}}^{2}}a_{y}+\frac{2\Delta}{D}a_{s}a_{y}+\frac{2\Delta}{D}\text{Re}a_{y}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}\;.$ (72)
@@
$=\frac{2|\omega|}{D}a_{1}+\frac{1}{l_{\text{so}}^{2}}a_{1}+\frac{2\Delta}{D}a_{s}a_{1}+\frac{n}{r}\gamma_{o}a_{s}\;,$ (73)
@@
$=\frac{2|\omega|}{D}a_{2}+\frac{1}{l_{\text{so}}^{2}}a_{2}+\frac{2\Delta}{D}(1+a_{s})a_{2}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}\;.$ (74)
@@
We confirm that for $n=0$ nonetheless $a_{s}$ and $a_{2}$ are the only non-vanishing terms, and hence the density of state has no spin-dependence. For $n\neq 0$ this is not the case and an excess spin is generated.
@@
The results are shown in Fig. 3, of which the left panel corresponds to Fig. 1 in the main text. We show the current density $I$ normalized by the maximum current density and the excess spin density $S^{l}$ normalized by the maximum, as a function of the radial coordinate, which is normalized by $\xi=\sqrt{D/2\Delta(T)}$. Clearly, for low temperatures, the spatial dependence of the excess spin density differs from the spatial dependence of the current density, with the spin density profile having its maximum further away from the vortex core. This is in contrast with the GL theory prediction, in which the spin depends linearly on the local current. For larger temperatures the difference becomes smaller. However, we notice that even for $T/T_{c}=0.98$, i.e. close to the critical current, there is still a noticeable difference between the two spatial profiles.
@@
Appendix E The Ginzburg-Landau limit: Derivation of the Lifshtiz-invariant
@@
$g\approx\begin{bmatrix}1-\frac{\Delta\tilde{\Delta}}{2}f_{0}\tilde{f}_{0}&\Delta f_{0}+\Delta^{2}\tilde{\Delta}f_{1}\\ \tilde{\Delta}\tilde{f}_{0}+\Delta\tilde{\Delta}^{2}\tilde{f}_{1}&-1+\frac{\Delta\tilde{\Delta}}{2}\tilde{f}_{0}f_{0}\end{bmatrix},$ (75)
where $f_{0,1}$ and $\tilde{f}_{0,1}$ are functions independent of $\Delta,\tilde{\Delta}$, which in real space are related by $\tilde{f}_{0,1}=f_{0,1}^{*}$. In this limit, only the derivative term, the energy term, the exchange field, the spin-galvanic term and the spin relaxation term remain in the equation. Indeed, we apply the Fourier transform and obtain
$(|\omega|+D_{ij}q_{i}q_{j})f_{0}-\Delta+i\text{sign}(\omega)(h_{k}+q_{l}\gamma_{kl})\{\sigma_{k},f_{0}\}-\Gamma_{ij}(f_{0}-\sigma_{i}f_{0}\sigma_{j})=0\;.$ (76)
From this equation it is clear that the spin-galvanic term appears only as an effective exchange field $q_{l}\gamma_{kl}$. In other words, it renormalizes the latter as $\tilde{h}_{k}=h_{k}+q_{l}\gamma_{kl}$. The term in the free energy due to an exchange or Zeeman field can be written as
$F_{Z}=\frac{1}{2}\chi_{ij}\tilde{h}_{i}\tilde{h}_{j}\;,$ (77)
@@
$F_{L}=\chi_{ij}h_{j}\gamma_{ik}\partial_{k}\varphi\;.$ (78)
This expression is identical to the expression for the Lifshitz invariant of the Ginzburg-Landau functional presented in [49] we define the second rank pseudotensor as $K_{jk}=\chi_{ij}\gamma_{ik}/\Delta^{2}$. However, since a phase gradient appears in the same way as a magnetic field even in the nonlinear equation, is the lowest order response to $h_{k}+q_{l}\gamma_{kl}$ for any temperature has the form of Eq. (77), with suitable definition of $\chi_{ij}(T)$. Therefore, Eq. (78) is more generic than the expression for the Lifshitz invariant in [49], it represents the contribution to the Free energy linear in the phase gradient for any temperature.

### v3

@@ Appendix E Vortex in a gyrotropic superconductor
Now, according to Eq. (9) in the main text, a gradient in the Green’s function is coupled to the spin via the term in the torque $\gamma_{kj}\sigma_{j}\partial_{k}g_{s}$, where $g_{s}$ is the Green’s function in the absence of a spin-galvanic effect. We may distinguish two different kinds of spatial dependence, spatial dependence of the phase and spatial dependence of the magnitude. Indeed, for a general system, we may write
$g_{s}(\bm{r})=e^{-i\phi(\bm{r})\tau_{3}/2}(g_{s0}(\bm{r})\tau_{3}+f_{s0}(\bm{r})\tau_{2})e^{i\phi(\bm{r})\tau_{3}/2}\;,$ (S91)
@@
$\bm{h}_{\text{eff},j}=n\bm{\hat{\phi}}_{k}\gamma_{kj}\;.$ (S92)
To derive the symmetry of the spin for generic forms of the spin-galvanic tensor, it is instructive to consider three cases separately, in all of which the tensor has two nonvanishing elements. First of all, we consider $\gamma_{xy}=-\gamma_{yx}$. A well-known example of spin-orbit coupling that leads to a spin-galvanic tensor of this form is Rashba spin orbit coupling. Generally it is the only allowed symmetry in the $C_{3v},C_{4v}$ and $C_{6v}$ point groups. For this symmetry the spin is always perpendicular to the direction of the current and hence around a vortex the spin points in the radial direction. The sign of $n\gamma_{xy}$ determines whether this is radially outward ($>0$) or inward ($<0$). Moreover, the magnitude of the excess spin is homogeneous at a fixed distance from the vortex core, it is independent of $\phi$.
@@
We now compute explicitly the current and excess spin densities by solving the Usadel equation, Eq. (7-9) in the main text with $\alpha_{kj}=0$, which read
$\partial_{k}\mathcal{J}_{k}+[\tau_{3}(\hat{\omega}+i\mathbf{h}{\bm{\sigma}})+\hat{\Delta},g]=\mathcal{T}-\frac{1}{8}\Gamma_{jk}[\sigma_{j}g\sigma_{k},g]\;,$ (S93)
$\mathcal{J}_{k}$ $=-Dg\tilde{\nabla}_{k}g+\frac{i}{16}\varepsilon_{ijl}\gamma_{lk}\{[\sigma_{i},g],\sigma_{j}+g\sigma_{j}g\}\;,$ (S94)
$\mathcal{T}$ $=-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]\;.$ (S95)
We assume an isotropic spin-relaxation $\Gamma_{jk}=\frac{D}{l_{\text{so}^{2}}}\delta_{jk}$, where $l_{\text{so}}$ is the spin-relaxation length, for simplicity of notation. For this we use the Riccati parameterization (we use the notation $a,b$ instead of the usual $\gamma,\tilde{\gamma}$ to avoid confusion with the tensor of the gyrotropic term)
$g=\begin{bmatrix}(1+ab)^{-1}(1-ab)\text{sign}(\omega_{n})&2(1+ab)^{-1}a\\ 2(1+ba)^{-1}b&-(1+ba)^{-1}(1-ba)\text{sign}(\omega_{n})\end{bmatrix}\;.$ (S96)
This parametrization guarantees the normalization condition $g^{2}=1$. For a single vortex, in the absence of a gyrotropic tensor ($\gamma_{lk}=0$) the equation for $a$ can be written as
$\nabla^{2}a-a\nabla a\cdot\nabla b=2\frac{|\omega|}{D}a+\frac{2\Delta}{D}e^{in\phi}(1-a^{2}-(be^{in\phi}-ae^{-in\phi}))+\frac{1}{l_{\text{so}}^{2}}(a-\sigma_{k}a\sigma_{k})\;,$ (S97)
@@
$\frac{\partial^{2}a}{\partial r^{2}}+\frac{1}{r}\frac{\partial a}{\partial r}+\frac{1}{r^{2}}\frac{\partial^{2}a}{\partial\phi^{2}}-a\Big{(}\frac{\partial a}{\partial r}\frac{\partial b}{\partial r}+\frac{1}{r^{2}}\frac{\partial a}{\partial\phi}\frac{\partial b}{\partial\phi}\Big{)}=\frac{2|\omega|}{D}a-\frac{\Delta}{D}(1-a^{2}-(be^{in\phi}-ae^{-in\phi}))+\frac{1}{l_{\text{so}}^{2}}(a-\sigma_{k}a\sigma_{k})\;.$ (S98)
@@
$a_{0}(r,\phi)=a_{s}(r)e^{in\phi}\;,$ (S99)
$b_{0}(r,\phi)=a_{s}(r)e^{-in\phi}\;.$ (S100)
The equation for $a_{s}(r)$ is a second order nonlinear equation with nonconstant coefficients in one variable;
$\frac{\partial^{2}a_{s}}{\partial r^{2}}+\frac{1}{r}\frac{\partial a_{s}}{\partial r}-\frac{n^{2}}{r^{2}}a_{s}-a_{s}\Big{(}\left(\frac{\partial a_{s}}{\partial r}\right)^{2}+\frac{n^{2}}{r^{2}}a_{s}^{2}\Big{)}=\frac{2|\omega|}{D}a_{s}-\frac{\Delta}{D}(1-a_{s}^{2})\;.$ (S101)
@@
$I(r)=\frac{-i}{1+a_{0}b_{0}}(a_{0}\nabla b_{0}-b_{0}\nabla a_{0})=\frac{1}{r}\frac{na_{s}^{2}}{1+a_{s}^{2}}\bm{\hat{\phi}}\;.$ (S102)
We note that while there exists a factor $\frac{1}{r}$ in the expression, the current density does not diverge near the center, since $a_{s}$ has to be at least first order in $r$ following Eq. (S101). Hence, the current density vanishes at $r=0$.
@@
$=\frac{2|\omega|}{D}\delta a_{t}+\frac{2\Delta}{D}a_{s}\delta b_{t}+\frac{\Delta}{D}(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi})+\gamma_{o}(\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}-\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{in\phi}\;,$ (S103)
@@
$=\frac{2|\omega|}{D}\delta b_{t}+\frac{\Delta}{D}(2a_{s}\delta a_{t}-(\delta b_{t}e^{in\phi}-\delta a_{t}e^{-in\phi}))-\gamma_{o}(-\frac{in}{r}a_{s}\sigma_{x}e^{i\phi\sigma_{z}}+\frac{\partial a_{s}}{\partial r}\sigma_{y}e^{i\phi\sigma_{z}})e^{-in\phi}\;.$ (S104)
@@
Again exploiting the symmetry of the problem we may write $\delta a_{t}(r,\phi)=(a_{x}(r)\sigma_{x}+a_{y}(r)\sigma_{y})e^{i\phi\sigma_{z}}e^{in\phi}$ and $\delta b_{t}(r,\phi)=\sigma_{y}\delta a_{t}(r,\phi)^{*}\sigma_{y}=-(a_{x}(r)^{*}\sigma_{x}+a_{y}(r)^{*}\sigma_{y})e^{i\phi\sigma_{z}}e^{-in\phi}$ to obtain two coupled nonlinear second order differential equations. We first consider the case in which the tensor of the gyrotropic term has the same symmetry as for Rashba spin orbit coupling, i.e. $\gamma_{xy}=-\gamma_{yx}=\gamma_{o}$, while all other terms of the tensor vanish. In this case
@@
$=\frac{2|\omega|}{D}a_{x}+\frac{1}{l_{\text{so}}^{2}}a_{x}+\frac{2\Delta}{D}a_{s}a_{x}+\frac{2\Delta}{D}\text{Re}a_{x}+\frac{in}{r}\gamma_{o}a_{s}\;,$ (S105)
@@
$=\frac{2|\omega|}{D}a_{y}+\frac{1}{l_{\text{so}}^{2}}a_{y}+\frac{2\Delta}{D}a_{s}a_{y}+\frac{2\Delta}{D}\text{Re}a_{y}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}\;.$ (S106)
@@
$=\frac{2|\omega|}{D}a_{1}+\frac{1}{l_{\text{so}}^{2}}a_{1}+\frac{2\Delta}{D}a_{s}a_{1}+\frac{n}{r}\gamma_{o}a_{s}\;,$ (S107)
@@
$=\frac{2|\omega|}{D}a_{2}+\frac{1}{l_{\text{so}}^{2}}a_{2}+\frac{2\Delta}{D}(1+a_{s})a_{2}-\gamma_{o}\frac{\partial a_{s}}{\partial_{r}}\;.$ (S108)
@@
We confirm that for $n=0$ nonetheless $a_{s}$ and $a_{2}$ are the only nonvanishing terms, and hence the density of state has no spin-dependence. For $n\neq 0$ this is not the case and an excess spin is generated.
@@
The results are shown in Fig. 3, of which the left panel corresponds to the figure in the main text. We show the current density $I$ normalized by the maximum current density and the excess spin density $S^{l}$ normalized by the maximum, as a function of the radial coordinate, which is normalized by $\xi=\sqrt{D/2\Delta(T)}$. Clearly, for low temperatures, the spatial dependence of the excess spin density differs from the spatial dependence of the current density, with the spin density profile having its maximum further away from the vortex core. This is in contrast with the GL theory prediction, in which the spin depends linearly on the local current. For larger temperatures the difference becomes smaller. However, we notice that even for $T/T_{c}=0.98$, i.e. close to the critical current, there is still a noticeable difference between the two spatial profiles.
@@
Appendix F The Ginzburg-Landau limit: Derivation of the Lifshtiz-invariant
@@
$g\approx\begin{bmatrix}1-\frac{\Delta\tilde{\Delta}}{2}f_{0}\tilde{f}_{0}&\Delta f_{0}+\Delta^{2}\tilde{\Delta}f_{1}\\ \tilde{\Delta}\tilde{f}_{0}+\Delta\tilde{\Delta}^{2}\tilde{f}_{1}&-1+\frac{\Delta\tilde{\Delta}}{2}\tilde{f}_{0}f_{0}\end{bmatrix}\;,$ (S109)
where $f_{0,1}$ and $\tilde{f}_{0,1}$ are functions independent of $\Delta,\tilde{\Delta}$, which in real space are related by $\tilde{f}_{0,1}=f_{0,1}^{*}$. In this limit, only the derivative term, the energy term, the exchange field, the spin-galvanic term, and the spin relaxation term remain in the equation. Indeed, we apply the Fourier transform and obtain
$(|\omega|+D_{ij}q_{i}q_{j})f_{0}-\Delta+i\text{sign}(\omega)(h_{k}+q_{l}\gamma_{kl})\{\sigma_{k},f_{0}\}-\Gamma_{ij}(f_{0}-\sigma_{i}f_{0}\sigma_{j})=0\;.$ (S110)
From this equation, it is clear that the spin-galvanic term appears only as an effective exchange field $q_{l}\gamma_{kl}$. In other words, it renormalizes the latter as $\tilde{h}_{k}=h_{k}+q_{l}\gamma_{kl}$. The term in the free energy due to an exchange or Zeeman field can be written as
$F_{Z}=\frac{1}{2}\chi_{ij}\tilde{h}_{i}\tilde{h}_{j}\;,$ (S111)
@@
$F_{L}=\chi_{ij}h_{j}\gamma_{ik}\partial_{k}\varphi\;.$ (S112)
This expression is identical to the expression for the Lifshitz invariant of the Ginzburg-Landau functional presented in [69] if we define the second rank pseudotensor as $K_{jk}=\chi_{ij}\gamma_{ik}/\Delta^{2}$. However, since a phase gradient appears in the same way as a magnetic field even in the nonlinear equation, the lowest order response to $h_{k}+q_{l}\gamma_{kl}$ for any temperature has the form of Eq. (S111), with suitable definition of $\chi_{ij}(T)$. Therefore, Eq. (S112) is more generic than the expression for the Lifshitz invariant in [69], it represents the contribution to the Free energy linear in the phase gradient for any temperature.